the perimeter of a rectangle is 234 meters. The width is 3 meters less than the length. Find the lenght and width.
Answer by BrittanyM(80)

We've been given the perimeter and the width in terms of legnth of a rectangle. With this we are asked to find both the legnth and width.

Perimeter = 234
W = L-3

And we know that the equation for perimeter is

P = 2L + 2W

So now we can plug in our given values to solve

234 = 2(L-3) + 2(L)

234 = 4L - 6
L = 60

Now, with this we can find the value of W.

W = L-3
W = 60-3
W = 57